OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is 6, the last term is 24,576, and the common ratio is −4? −32,766 −19,662 19,662 32,766

OpenStudy (anonymous):

847

OpenStudy (doc.brown):

First number = 6 Second number = 6*(-4) = -24 Third number = -24*(-4) = 96 etc

OpenStudy (anonymous):

Sum of a geometric series is given by: \[\huge S_n= \frac{a(r^n -1)}{r-1}\] \[\huge S_7= \frac{6[(-4)^7 -1]}{r-1}\] \[\huge S_7= \frac{6[-16384 -1]}{-4-1}\] \[\huge S_7= \frac{6[-16385]}{-5}\] Now solve it....